OpenStudy (anonymous):

Find tan θ if sec θ = (sqrt. 37)/(6) and sin θ < 0

OpenStudy (anonymous):

answer is 1/6 haci

OpenStudy (anonymous):

 that's what I have set up so far for tangent... mind explaining?

OpenStudy (anonymous):

how 'bout using the fact that tg^2(theta)+1=sec^2(theta)

OpenStudy (anonymous):

tg^2(theta)+1=37/36 /add (-1) tg^2(theta)=1/36 tg(theta)=+-1/6 I guess you can finish it